OpenStudy (anonymous):

Show all steps used to solve the problem. One point for steps and one point for solution. Be sure that you check your solution for correctness. 17-x/4 =-10

OpenStudy (anonymous):

7 -1x/4 = -10 -1x/4 = -10 -7 -1x/4 = -17 (-4/1)(-1x/4) = (-17)(-4/1) x =(-17)(-4)

OpenStudy (anonymous):

thanks